OpenStudy (anonymous):

Solving Exponential Equations with Unequal Bases 2x + 3 = 7

OpenStudy (turingtest):

what is all this about exponential equations...?

OpenStudy (anonymous):

Did you read what Turing posted in your earlier question?

OpenStudy (anonymous):

I need help, obviously.

OpenStudy (turingtest):

But I am afraid you did not type the question properly. It's like saying "how tall is this building?" and then you show me a tree...

OpenStudy (anonymous):

no let me re-write the problem.

OpenStudy (turingtest):

if the questions is just 2x + 3 = 7 then it is a simple linear equation

OpenStudy (turingtest):

ah...

OpenStudy (turingtest):

you should have written this as 2^(x+3)=7 take the log base 2 of both sides, then tell me wht you have

OpenStudy (anonymous):

\( x = \log_2 7 -3 \)

OpenStudy (anonymous):

2^(x+3)=7 log 2^(x+3)= log 7 (x + 3) = log(2)7 x = log(2)7 -3

OpenStudy (anonymous):

or , x = log7/ log2 -3

OpenStudy (anonymous):

Sorry, but im not comprehending any of this.

OpenStudy (anonymous):

which part

OpenStudy (anonymous):